Calculate the energy density in an electric field of 106 V/m


1
Expert's answer
2021-05-18T10:56:06-0400

Given,


"E = 10^6V\/m"

We know that,


"\epsilon_0 = 8.85 \times 10^{-12}"

The energy density formula of the capacitor is given by:


"U = \dfrac{1}{2}\epsilon_0E^2"

"U = \dfrac{1}{2}(8.85 \times 10^{-12} \times 10^{12})"

"U = 4.425"